OpenStudy (anonymous):

Please help! If a total of 679 adult tickets were sold for a school play. There were 71 fewer student tickets sold than adult tickets. How many adult tickets were sold?

OpenStudy (tyrinzoebrown):

I wuld say 608 tickets

OpenStudy (anonymous):

it has to be more than just subtracting 71 from 679...

OpenStudy (anonymous):

Since a total of 679 tickets were sold, we can use this equation: s + a = 679 Then, there were 71 fewer students sold, that means more adult tickets were sold. s=a-71 Now put these in a system of equations and solve, firstly plug in "s=a-71" s + a = 679 s=a-71

OpenStudy (anonymous):

a^2 - 71a = 679?

OpenStudy (anonymous):

I would suggest substitution: (a-71) + a = 679 2a-71=679 2a = 750 a = 375

OpenStudy (anonymous):

xD nicely put there bluxar

OpenStudy (anonymous):

oh okay, I accidently distributed. But thanks for the help! I appreciate it!